• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Armstrong

    Log entry time 19:20:39 on October 24, 2009

    Entry number 296469

    keyword=Heidi saves the day...

    So, Heidi F. had called and, having read the elog, suggested that we try increasing the lead flows on Right Q2 by 10% or so. She describes the what and where. The idea being that the flow meters may have lost their correct calibrations, and the actual flows may be lower than what is read out...

    Initially, we read 90 l/m on the left readout and 120 l/m on the right readout from Hall A Tools (I presume that the right and left on the Tools screen are the same as the right and left labels on the valves).

    So, Wouter and I go to the Hall. The valves are in a roped-off area, due to radiation, so we call MCC to have an ARM come over. The ARM checks, and the area is "cool" now (it has been a while since we had beam!), and so we can get to the valves. I move each valve by about 1 mm counterclockwise; Alex is on the phone in the counting house, watching flows. Flows go up to 104 l/m and 135 l/m.

    Now we turn on the magnet. Ramps happily all the way to 1400 A!

    Now set to correct current for this momentum. All magnets on LHRS are cycled, at correct currents and regulating.

    Exit Hall, and go to Beam Permit, with fingers crossed.